OpenStudy (maheshmeghwal9):

Using Euclid's Division Lemma find the H.C.F. of 135 & 225.

Parth (parthkohli):

I don't exactly know that, but I still can find it. Would that be fine by you?

OpenStudy (maheshmeghwal9):

np:)

Parth (parthkohli):

Okay, first find the factors of both and reply that to me :D

OpenStudy (maheshmeghwal9):

k! wait please.

OpenStudy (maheshmeghwal9):

135=3*3*3*5 & 225=3*3*5*5

Parth (parthkohli):

Take the common factors of both and then multiply.

OpenStudy (maheshmeghwal9):

3*3*5=45,then

Parth (parthkohli):

45 is the HCF :P

Parth (parthkohli):

I don't know that Euclid's lemma or whatever

OpenStudy (maheshmeghwal9):

k!dude..... Thanx!.:)